please help me solve this word problem .....there are two consecutive numbers.the sum of the lesser number and three times the greater number is is 103.what are the two numbers
Answer by flame8855(424)

x
x+1
x + 3(x+1) =103
x + 3x+3 =103
4x = 100
x = 25
x+1 = 26